# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ 2.3.4 Beyond Sequence Patterns
To deal with sets, we can use the same tuple pattern expression as given in Section 2.1, and interpret the concatenation as the (disjoint) set union operator. We identify patterns $p$ up to the permutation and introduce the following rewrite rules.
@@
Let $M$ be: $\left(\begin{array}[]{rrr}\{a\}&\ \{b\}&\ \{a,b\}\\ \{b\}&\ \{b,c\}&\{b,c\}\end{array}\right)$. The pattern $(xz,yz,xyz)$ (which means that the third set is a superset of the first and second sets, where $z$ is the intersection of the first and second sets) is inferred as follows.
$((x_{1},x_{2},x_{3}),[M/(x_{1},x_{2},x_{3})])$ $\longrightarrow((x_{1},x_{2},x_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime}\\ \{a\}&\ \{b\}&\ \{b\}\\ \{b\}&\ \{b,c\}&\{c\}\end{array}\right))$ (R-Subset)
$\longrightarrow((x_{1},x_{3}^{\prime}x^{\prime}_{2},x_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}^{\prime}&x_{3}^{\prime}\\ \{a\}&\ \emptyset&\ \{b\}\\ \{b\}&\ \{b\}&\{c\}\end{array}\right))$ (R-Subset)
$\longrightarrow((x_{2}^{\prime}x^{\prime}_{1},x_{3}^{\prime}x^{\prime}_{2},x_{2}^{\prime}x^{\prime}_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x^{\prime}_{1}&x_{2}^{\prime}&x_{3}^{\prime}\\ \{a\}&\ \emptyset&\ \{b\}\\ \emptyset&\ \{b\}&\{c\}\end{array}\right))$ (R-Subset)
We can deal with multisets in a similar manner, where $xy$ is interpreted as the multiset union of $x$ and $y$.
3 Applications to Automated Verification of List-Manipulating Programs
In this section, we discuss how to apply the STP inference to automated verification of programs manipulating list-like data structures. We focus on CHC-based automated program verification [7], where program verification problems are reduced to the satisfiability problem for constrained Horn clauses (CHCs, a.k.a. constraint logic programs), and passed to an automated CHC solver. The CHC-based approach provides a uniform, language-agnostic approach to automated verification. It has recently been studied extensively and applied to various languages, ranging from low-level languages like C and LLVM [22, 51], to high-level languages like Java [29], Rust [37], and OCaml [52, 8].
We first review the definition of CHCs and their connection with program verification in Section 3.1. We then propose a basic application of STP inference to CHC solving over list-like data structures in Section 3.2. Sections 3.3 and 3.4 discuss further enhancement of the STP-based CHC solving proposed in Section 3.2.
3.1 Constrained Horn Clauses
@@
It is satisfiable. Indeed, $\{\mathit{Plus}\mapsto\{(x,y,z)\mid x\geq 0\land z=x+y\}\}$ satisfies all the three clauses. Henceforth, we often use a formula to express a model, like $\mathit{Plus}(x,y,z)\equiv x\geq 0\land z=x+y$. A model is not unique in general; for the example above, $\mathit{Plus}(x,y,z)\equiv z\geq y$ is also a model. The CHC satisfiability problem is undecidable in general, but a number of CHC solvers, which automatically solve the problem have been constructed (though they are inevitably incomplete). They typically support integer/real arithmetic, arrays, and ADTs. A competition on CHC solvers called CHC-COMP is held annually: see https://chc-comp.github.io/.
Various program verification problems can be reduced to the CHC problem. For example, the system of CHCs above has been obtained from the assertion safety of the following program:
let rec plus x y = if x=0 then y else (plus (x-1) y)+1
let main x y = assert(plus x y >= y).
The predicate $\mathit{Plus}$ denotes the invariant between the arguments $x,y$ and the return value $z$ of plus. The system of CHCs is satisfiable if and only if $\texttt{main}\;x\;y$ does not suffer from assertion failures for any integers $x$ and $y$.
From a program that uses ADTs, we can obtain a system of CHCs over ADTs, as in the reva example in Section 1. Many of the CHC solvers [24, 32, 9, 53] support ADTs, but they are not strong enough; improving CHC solvers over ADTs has been the main motivation for the present work.
Below we often make use of the following well-known fact. For completeness we show a proof.
Fact 3.1.
Let $S=D\cup G$ be a system of CHCs, where $D$ and $G$ consist of definite and goal clauses respectively, and all the predicates occurring in $G$ also occur in $D$. Then, there exists a least model of $D$, and the followings are equivalent.
  1. 1.
    $S$ is satisfiable.
  2. 2.
    The least model of $D$ satisfies $G$.
Furthermore, suppose $G^{\prime}$ is another set of goal clauses. If the least model of $D$ satisfies $G^{\prime}$, then $S$ is satisfiable if and only if $S\cup G^{\prime}$ is satisfiable.
@@
The existence of the least model of $D$ follows from Knaster-Tarski fixpoint theorem (where two models are ordered by the component-wise subset relation: $\theta_{1}\leq\theta_{2}$ just if $\theta_{1}(P)\subseteq\theta_{2}(P)$ for every predicate variable $P$). It is obvious that (2) implies (1). To show the converse, let $\theta_{0}$ be the least model of $D$ and suppose $\theta$ satisfies $S$. Since $\theta$ is also a model of $D$, $\theta_{0}\leq\theta$. Since $\theta$ is a model of $G$ and predicate variables occur only negatively in $G$, $\theta_{0}$ is also a model of $G$. Thus (1) implies (2). The remaining claim is just a corollary. Suppose the least model of $D$ satisfies $G^{\prime}$, and $S=D\cup G$ is satisfiable. Then, by (2), the least model of $D$ also satisfies $G$, and hence also satisfies $S\cup G^{\prime}=D\cup G\cup G^{\prime}$. The converse is trivial. ∎
3.2 Direct Applications of STP Inference
We now discuss how to apply STP inference to CHC solving. For the sake of simplicity, we focus on solving CHCs over integer lists and integers, such as the one given in Section 1.
Before showing a concrete procedure, let us go through a running example.
@@ 3.2 Direct Applications of STP Inference
Recall the $\mathit{Reva}$ example in Section 1.
[…]
Here, we have identified lists with sequences. Let $D$ be the set consisting of the definite clauses (2) and (3), and $G$ be the singleton set consisting of the goal clause (1).
We first collect positive samples from $D$. From the clause (2), we know $\mathit{Reva}(\epsilon,0,0)$ and $\mathit{Reva}(\epsilon,1\cdot 2,1\cdot 2)$ must hold, for example. (Recall that all the first-order variables are implicitly universally quantified.) Thus, let the initial learning data $M$ be: $\left(\begin{array}[]{rrr}\epsilon&0&0\\ \epsilon&1\cdot 2&1\cdot 2\end{array}\right)$. By invoking the STP inference algorithm, we obtain the STP $(\epsilon,l,l)$. Thus, we obtain a candidate model $\mathit{Reva}(l_{1},l_{2},l_{3})\equiv l_{1}=\epsilon\land l_{2}=l_{3}$.
Next, we check whether the candidate model is indeed a model of $D$. To this end, we substitute the candidate model for the clauses and check the validity of the resulting formulas by using an SMT solver. From the clause (2), we obtain: $\forall l_{1},l_{2},l_{3}.\epsilon=\epsilon\land l_{2}=l_{2}$, which is obviously valid. From the clause (3), however, we obtain:
$\forall x,l^{\prime}_{1},l_{2},l_{3}.l_{1}^{\prime}=\epsilon\land x\cdot l_{2}=l_{3}\Rightarrow x\cdot l^{\prime}_{1}=\epsilon\land l_{2}=l_{3},$
which is invalid. Suppose that an SMT solver produces $x=0,l_{1}^{\prime}=l_{2}=\epsilon,l_{3}=x$ as a counterexample. Then, we add $(x\cdot l_{1}^{\prime},l_{2},l_{3})=(0,\epsilon,0)$ to the learning data $M$.
By re-running the STP inference algorithm, we may obtain an STP $(l_{1},l_{2},l_{1}l_{2})$, i.e., a candidate model $\mathit{Reva}(l_{1},l_{2},l_{3})\equiv l_{3}=l_{1}l_{2}$. Again, it does not satisfy the clause (3). Assuming that an SMT solver generates a counterexample $x=0,l_{1}^{\prime}=1,l_{2}=\epsilon,l_{3}=1\cdot 0$, we obtain $(x\cdot l_{1}^{\prime},l_{2},l_{3})=(0\cdot 1,\epsilon,1\cdot 0)$ as a new sample. By performing STP inference again, we obtain a pattern $(l_{1},l_{2},l_{1}^{R}l_{2})$ and the corresponding candidate model $\mathit{Reva}(l_{1},l_{2},l_{3})\equiv l_{3}=l_{1}^{R}l_{2}$, which satisfies $D$.
We now proceed to check whether $\mathit{Reva}(l_{1},l_{2},l_{3})\equiv l_{3}=l_{1}^{R}l_{2}$ also satisfies the goal clause (1), i.e., whether the formula:
$\forall l_{1},l_{2},l_{3},l_{4},l_{5}.l_{3}=l_{1}^{R}l_{2}\land l_{4}=l_{3}^{R}\land l_{5}=l_{2}^{R}l_{1}\Rightarrow l_{4}=l_{5}.$
is valid. CVC5, a state-of-the-art SMT solvers supporting the sequence theory, can instantly answer yes. Thus, we have checked that the system of CHCs $D\cup G$ is satisfiable. ∎
1:function solve($D,G$)
2:  $T$ := $\emptyset$;
3:  $M$ := $\mathtt{collect\_samples}(D)$;
4:  while true do
5:   if $\mathtt{true\_samples}(M)\models\neg G$ then return UNSAT;
6:   end if
7:   $t_{\mathit{new}}$ := $\mathtt{STPinf}(M)$;
8:   if $t_{\mathit{new}}\not\in T$ and $T\cup\{t_{\mathit{new}}\}\models D$ then
9:     $T$ := $T\cup\{t_{\mathit{new}}\}$;
10:     if $T\models G$ then return SAT($T$)
11:     end if;
12:   else
13:     $M$ := $M$ $\cup\mathtt{collect\_more\_samples}(D,t_{\mathit{new}})$;
14:   end if
15:  end while
16:end function
Procedure 1 A Procedure for Solving CHCs Using STP Inference
We now explain a general procedure for solving CHCs by using STP inference, shown in Figure 1. The procedure solve takes a set $D$ of definite clauses and a set $G$ of goal clauses, and checks whether $S=D\cup G$ is satisfiable. For simplicity, we assume that $S$ contains a single predicate symbol $P$; an extension to deal with multiple predicates is straightforward. We write $m_{D,P}$ for the least model of $D$, i.e., the least relation such that $\{P\mapsto m_{D,P}\}$ satisfies $D$. We keep in $T$ a set of valid tuple patterns satisfied by the least model of $D$; in other words $T$ is a set of invariants on $P$ found so far. It is initially an empty set (line 2). On line 3, $\mathtt{collect\_samples}(D)$ returns a (finite) set of positive samples, i.e., a finite subset of $m_{D,P}$. For that, we can either apply unit propagations (i.e., resolutions where one of the two clauses is an atomic formula P(v~)P($\widetilde${v})) to $D$, or (if a source program is available) randomly execute the program to collect values $\widetilde{v}$ such that P(v~)P($\widetilde${v}) holds.
In the loop beginning at line 4, we expand $M$ and $T$ until the satisfiability of $S$ is proved or disproved. On line 5, $\mathtt{true\_samples}(M)$ filters out samples that may not belong to $m_{D,P}$ (such samples may have been introduced on line 13) and returns a subset $M^{\prime}\subseteq m_{D,P}$. We then check whether $\{P\mapsto M^{\prime}\}$ is a model of $\neg G$ (which is expressed as $M^{\prime}\models\neg G$ by abuse of notation); if so, $m_{D,P}\models\neg G$ also holds, since $P$ occurs only positively in $\neg G$ and $\mathtt{true\_samples}(M)\subseteq m_{D,P}$; hence we conclude that $S$ is unsatisfiable. [^8] Since this method for proving unsatisfiability would not be efficient, in practice, we run another procedure for proving the unsatisfiability in parallel, as reported in Section 4.
On line 7, we use the STP inference algorithm to obtain an STP $t_{\mathit{new}}$ such that $M\subseteq\mathcal{L}(t_{\mathit{new}})$. The subprocedure $\mathtt{STPinf}{}$ runs the algorithm in Section 2 in a non-deterministic manner.
Line 8 checks whether $t_{\mathit{new}}$ is a new pattern that satisfies $D$. Here, $T^{\prime}\models D$ means that $\{P\mapsto\bigcap_{t\in T^{\prime}}\mathcal{L}(t)\}$ is a model of $D$. That condition can be checked by using an SMT solver as follows. For $t=(p_{1},\ldots,p_{k})$ with variables $\widetilde{z}$, let $\varphi_{t}$ be $\lambda(x_{1},\ldots,x_{k}).\exists\widetilde{z}.(x_{1}=p_{1}\land\cdots x_{k}=p_{k})$ (where $\{x_{1},\ldots,x_{k}\}\cap\{\widetilde{z}\}$$\emptyset$). Then, the condition $(x_{1},\ldots,x_{k})\in\mathcal{L}(t)$ is equivalent to $\varphi_{t}(x_{1},\ldots,x_{k})$. Thus, $\{P\mapsto\bigcap_{t\in T^{\prime}}\mathcal{L}(t)\}$ satisfies $D$, if and only if $\lambda\widetilde{x}.\bigwedge_{t\in T^{\prime}}\varphi_{t}/P]D$ is valid. The latter formula is a first-order formula consisting of string constants and operations; hence its validity can be checked by an SMT solver supporting sequences.

### v2

@@ 4.2 Beyond Sequence Patterns
To deal with multisets, we just need to interpret the empty pattern and the pattern concatenation as the emptyset and the multiset union, respectively, and identify patterns up to permutations (for example, $xyz=yzx$). Inferences rules for tuple patterns and reduction rules for patterns can be adjusted accordingly, as follows.
@@
$\frac{\begin{array}[]{@{}c@{}}p_{j}=p_{i}p^{\prime}_{j}\quad\quad p_{i}\neq\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-Subset)
$\frac{\begin{array}[]{@{}c@{}}p_{j}=ap^{\prime}_{j}\quad\quad a\in\Sigma\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-CSubset)
$\frac{\begin{array}[]{@{}c@{}}p_{j}=\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-Empty)
Most of the results for STPs and CSTPs carry over to their multiset version (written SMTPs and CSMTPs), including the soundness and completeness of the inference algorithm, Theorem 3.7, and Theorem 3.10 (where word equations are replaced with multiset equations). A notable exception is that the minimality property (Theorem 3.4) fails, but it is not an obstacle to the application to program verification discussed in the next section. Similarly, we can deal with sets by interpreting the concatenation as the disjoint union.
@@
Let $M$ be: $\left(\begin{array}[]{rrr}\{a,a\}&\ \{a,b\}&\ \{a,a,b\}\\ \{b\}&\ \{c\}&\{b,c\}\end{array}\right)$. The pattern $(xz,yz,xyz)$ (which means that the third set is a superset of the first and second sets, where $z$ is the intersection of the first and second sets) is inferred as follows.
$\Big((x_{1},x_{2},x_{3}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}\\ \{a,a\}&\ \{a,b\}&\ \{a,a,b\}\\ \{b\}&\ \{c\}&\{b,c\}\end{array}\right)\Big)\longrightarrow\Big((x_{1},x_{2},x_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime}\\ \{a,a\}&\ \{a,b\}&\ \{b\}\\ \{c\}&\ \{c\}&\{c\}\end{array}\right)\Big)$
$\longrightarrow\Big((x_{1},x_{3}^{\prime}x^{\prime}_{2},x_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}^{\prime}&x_{3}^{\prime}\\ \{a,a\}&\ \{a\}&\ \{b\}\\ \{c\}&\ \emptyset&\{c\}\end{array}\right)\Big)\longrightarrow\Big((x_{2}^{\prime}x^{\prime}_{1},x_{3}^{\prime}x^{\prime}_{2},x_{2}^{\prime}x^{\prime}_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x^{\prime}_{1}&x_{2}^{\prime}&x_{3}^{\prime}\\ \{a\}&\ \{a\}&\ \{b\}\\ \{c\}&\ \emptyset&\{c\}\end{array}\right)\Big).$
5 Applications to Automated Verification of
List-Manipulating Programs
This section discusses how to apply the (C)STP inference to automated verification of programs manipulating list-like data structures. We focus on CHC-based automated program verification [6], where program verification problems are reduced to the satisfiability problem for constrained Horn clauses (CHCs, a.k.a. constraint logic programs), and passed to an automated CHC solver. The CHC satisfiability problem is undecidable in general, but a number of CHC solvers that automatically solve it have been developed, though they are inevitably incomplete. These solvers typically support integer and real arithmetic, arrays, and ADTs. An annual competition on CHC solvers called CHC-COMP is held; see https://chc-comp.github.io/. The CHC-based approach provides a uniform, language-agnostic approach to automated verification. It has recently been studied extensively and applied to various languages, ranging from low-level languages like C and LLVM [23, 59], to high-level languages like Java [30], Rust [42], and OCaml [60, 7]. As noted in Section 1, however, CHC solvers over ADTs have so far been unsatisfactory.
Below we first introduce a class of CHCs with word constraints, and show that it is decidable whether a system of CHCs (i.e., a finite set of CHCs) in that class has a model describable using CSTPs in Section 5.1. This result is directly applicable to the $\mathit{Reva}{}$ example in Section 1. Sections 5.2 and 5.3 discuss further enhancement of the STP-based CHC solving.
5.1 Constrained Horn Clauses over Words
@@
It is satisfiable. Indeed, $\{\mathit{Plus}\mapsto\{(x,y,z)\mid x\geq 0\land z=x+y\}\}$ satisfies all the clauses. Henceforth, we often use a formula to express a model, like $\mathit{Plus}(x,y,z)\equiv x\geq 0\land z=x+y$. A model is not unique in general; for the example above, $\mathit{Plus}(x,y,z)\equiv z\geq y$ is also a model.
In this section, instead of CHCs over integer arithmetic, we consider CHCs on words over $\Sigma$. Thus, each variable ranges over the set $\Sigma^{*}$ of words, and constraints $\varphi$ are drawn from the set of quantifier-free STP formulas as defined in Definition 3.1.
Example 5.1.
The $\mathit{Reva}$ example in Section 1 is expressed as the following system of CHCs $\mathcal{S}_{\mathit{Reva}}$.
$\mathit{Reva}(\epsilon,l_{2},l_{2}).\qquad\mathit{Reva}(l^{\prime}_{1},xl_{2},l_{3})\land|x|=1\Rightarrow\mathit{Reva}(xl^{\prime}_{1},l_{2},l_{3}).$
$\mathit{Reva}(l_{1},l_{2},l_{3})\land\mathit{Reva}(l_{3},\epsilon,l_{4})\land\mathit{Reva}(l_{2},l_{1},l_{5})\Rightarrow l_{4}=l_{5}.$
Here, the constraint $|x|=1$ can be expressed by $\bigvee_{a\in\Sigma}x=a$, assuming that the alphabet $\Sigma$ is finite.
Let $t=(x,y,x^{R},y)$. Then $\{\mathit{Reva}\mapsto\mathcal{L}(x,y,x^{R}y)\}$ is a model of the CHCs above. ∎
We say that a system $\mathcal{S}$ of CHCs on words has a CSTP-model if there exist $\gamma_{1},\ldots,\gamma_{n}$ such that $\{P_{1}\mapsto\mathcal{L}(\gamma_{1}),\ldots,P_{n}\mapsto\mathcal{L}(\gamma_{n})\}$ is a model of $\mathcal{S}$. The following is the main result of this section.
Theorem 5.1.
Given a system $\mathcal{S}$ of CHCs on words, it is decidable whether $\mathcal{S}$ has a CSTP-model.
Proof Sketch..
Let $\mathcal{S}=\mathcal{S}_{D}\cup\mathcal{S}_{G}$, where $\mathcal{S}_{D}$ and $\mathcal{S}_{G}$ respectively consist of definite and goal clauses. By Theorem 3.7, the least CSTP-model $\mathcal{M}_{0}=\{P_{1}\mapsto\mathcal{L}(\gamma_{1}),\ldots,P_{n}\mapsto\mathcal{L}(\gamma_{n})\}$ of $\mathcal{S}_{D}$ exists and it is computable (note that the algorithm for $g_{\mathcal{F}}$ can be constructed by using Theorem 3.10). Then, $\mathcal{S}$ has a CSTP-model if and only if $\mathcal{M}_{0}$ is a model of $\mathcal{S}_{G}$. The latter is decidable by Theorem 3.10. ∎
An analogous result holds also when restricted to STP-models. We say that a system $\mathcal{S}$ of CHCs on words has a STP-model if there exist STPs $t_{1},\ldots,t_{n}$ such that $\{P_{1}\mapsto\mathcal{L}(t_{1}),\ldots,P_{n}\mapsto\mathcal{L}(t_{n})\}$ is a model of $\mathcal{S}$. Here, we add a special STP $t_{\emptyset}$ that denotes the empty set (i.e., $\mathcal{L}(t_{\emptyset})=\emptyset$).
Theorem 5.2.
Given a system $\mathcal{S}$ of CHCs on words, it is decidable whether $\mathcal{S}$ has an STP-model.
To prove the theorem, it suffices to use Theorem 3.8 instead of Theorem 3.7.
Proof Sketch..
Let $\mathcal{S}=\mathcal{S}_{D}\cup\mathcal{S}_{G}$, where $\mathcal{S}_{D}$ and $\mathcal{S}_{G}$ respectively consist of definite and goal clauses. By Theorems 3.8 and 3.10, we can enumerate all the minimal STP-models $\mathcal{M}_{0}=\{P_{1}\mapsto\mathcal{L}(t_{1}),\ldots,P_{n}\mapsto\mathcal{L}(t_{n})\}$ of $\mathcal{S}_{D}$. Then, it suffices to check whether one of the minimal STP-models is also a model of $\mathcal{S}_{G}$. ∎
The unsatisfiability of CHCs (where models are not restricted to CSTPs) is also semi-decidable:
Theorem 5.3.
There exists a procedure which, given a system $\mathcal{S}$ of CHCs on words, eventually outputs “Unsat” whenever $\mathcal{S}$ is not satisfiable.
@@
This follows from the fact that $\mathcal{S}$ is unsatisfiable if and only if there is a resolution proof. Thanks to the decidability of word equations, we just need to enumerate a candidate resolution proof, and check whether it is a valid proof. ∎
Note, however, that the satisfiability of CHCs is undecidable in general, when models are not restricted to CSTPs. To see this, notice that natural numbers $n$ can be encoded as the word $a^{n}$.
Below, we demonstrate how we can automatically obtain a CSTP-model for the $\mathit{Reva}$-example. For the sake of simplicity, we omit constant patterns.
@@
Recall the CHCs $\mathcal{S}_{\mathit{Reva}}$ in Example 5.1. We first compute the least model of the definite clauses (i.e., the first two clauses) $\mathcal{S}_{D}$, based on the algorithm in Theorem 3.7. First, we set $M_{0}=\emptyset$ and $\gamma_{0}=(a,\epsilon,\epsilon)\land(\epsilon,\epsilon,\epsilon)$, and check whether $\gamma_{0}$ satisfies $\mathcal{S}_{D}$. Using the algorithm in Theorem 3.10, we obtain an element $\widetilde{s}_{0}=(\epsilon,ab,ab)$ that should belong to the least model.
Then, we set $M_{1}=\{(\epsilon,ab,ab)\}$, and get $\gamma_{1}=\mathtt{CTPinf}(\{(\epsilon,ab,ab)\})=(\epsilon,x,x)$. This does not satisfy the second clause. Indeed, the formula obtained by negating the second clause: $(l^{\prime}_{1},xl_{2},l_{3})\in\mathcal{L}(\gamma_{1})\land|x|=1\land(xl^{\prime}_{1},l_{2},l_{3})\not\in\mathcal{L}(\gamma_{1})$ has a model $(x,l^{\prime}_{1},l_{2},l_{3})=(a,\epsilon,b,ab)$, from which we obtain a new element $\widetilde{s}_{1}=(a,b,ab)$. By setting $M_{2}=\{(\epsilon,ab,ab),(a,b,ab)\}$, we obtain $\gamma_{2}=(x,y,xy)\land(x,y,x^{R}y)$. Since $(l^{\prime}_{1},xl_{2},l_{3})\in\mathcal{L}(\gamma_{2})\land|x|=1\land(xl^{\prime}_{1},l_{2},l_{3})\not\in\mathcal{L}(\gamma_{2})$ has a model $(x,l^{\prime}_{1},l_{2},l_{3})=(a,b,\epsilon,ba)$, we obtain the new element: $\widetilde{s}_{2}=(ab,\epsilon,ba)$.
By setting $M_{3}=\{(\epsilon,a,a),(a,b,ab),(ab,\epsilon,ba)\}$, we obtain $\gamma_{3}=(x,y,x^{R}y)$. Now, $\gamma_{3}$ is a model of $\mathcal{S}_{D}$. Since $\gamma_{3}$ is also a model of the goal clause, we can conclude that $\mathcal{S}_{\mathit{Reva}}$ is satisfiable, and hence also that the reva program in Section 1 does not suffer from assertion failures. ∎
In Appendix B, we extend CSTP-models to piecewise CSTP-models, and apply them to verification of functional queues.